# A Comprehensive Survey on Continual Learning in Generative Models

### v3

@@ 3.1.1 Training objectives
$\epsilon_{i}\geq 0;\forall i\in[1,\cdots,t-1],$
where $\epsilon_{i}$ is a slack variable that permits a slight increase in loss on previous tasks, allowing controlled forgetting while prioritizing learning on the current task.
3.1.2 Backbone
In natural language generation tasks, mainstream LLMs generally fall into two architectural categories: encoder-decoder architectures and decoder-only architectures. According to recent literature, we summarize the backbone architectures commonly adopted in LLM’s continual learning task as follows:
  • BERT [167] employs a bidirectional encoder-only architecture based on the Transformer framework. Unlike autoregressive models, BERT leverages masked language modeling and next-sentence prediction as its pretraining objectives, enabling it to learn contextualized word representations from both left and right contexts simultaneously. Each encoder layer comprises multi-head self-attention, feedforward networks, layer normalization, and residual connections.
  • T5 [168] adopts a standard encoder-decoder framework, where both the encoder and decoder are built upon the Transformer architecture. T5 reformulates all NLP tasks into a unified text-to-text format. Each layer in T5 consists of conventional multi-head attention and feedforward sublayers, with configurable embedding dimensions and attention heads depending on the model scale.
  • GPT [169] series exemplifies a unidirectional decoder-only architecture, characterized by left-to-right autoregressive generation. GPT is composed of stacked decoder blocks, each containing masked multi-head attention, feedforward layers, and residual connections. The design is streamlined and restricts visibility to a leftward context, making it well-suited for open-ended generation tasks. In most research settings, the GPT-2 model is used.
  • LLaMA [13] series retains the decoder-only design while introducing several architectural optimizations, including Rotary Positional Embeddings (RoPE), the SwiGLU activation function, and RMSNorm. These improvements bolster training stability, accelerate inference, and preserve performance on tasks requiring extended context lengths.
3.1.3 Benchmark
Several benchmark frameworks exist to evaluate the continual learning performance of LLMs. In this part, we focus on some of the most representative benchmarks to provide a detailed introduction.
  • Short-sequence CL  [170] comprises five large-scale text classification datasets, including AG News (news classification, 4 classes), Amazon Reviews (sentiment analysis, 5 classes), Yelp Reviews (sentiment analysis, 5 classes), DBpedia (Wikipedia text classification, 14 classes), and Yahoo Answers (Q&A classification, 10 classes).
  • Long-sequence CL  [69] builds on the five datasets above by adding four tasks from the GLUE benchmark (MNLI, QQP, RTE, SST-2), five tasks from SuperGLUE (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB dataset, forming a longer continual learning evaluation sequence.
  • SuperNI [171] is a large-scale benchmark comprising 1,616 NLP tasks paired with natural language instructions, spanning 76 task types and 55 languages. Each task is accompanied by a clear task definition and illustrative examples demonstrating the mapping from input to the desired output. A representative continual learning workflow is often simulated using a diverse subset of tasks selected across domains and types from this benchmark.
  • TRACE [172] includes eight challenging datasets spanning domain-specific knowledge, multilingual ability, code generation, and mathematical reasoning. All datasets are uniformly formatted to allow automated evaluation. To measure forgetting in aligned LLMs during continual learning, TRACE introduces three core metrics: (1) General Ability Delta: tracks changes in performance on general tasks such as factual QA, reasoning, multilingual QA, commonsense reasoning, and reading comprehension; (2) Instruction Following Delta: measures shifts in instruction adherence on Self-instruct and LIMA prompt sets; (3) Safety Delta: evaluates changes in response safety using the CoNa dataset.
TABLE I: Summary of continual learning methods for Large Language Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Setting: CToD refers to Continual Task-oriented Dialogue, CIT to Continual Instruction Tuning, CPT to Continual Pre-Training, and CRE to Continual Relation Extraction.
Methods Taxonomy Setting Backbone / Basic models Implement Code
Arch Reg Rep
TPEM [60] CToD GLMP [173] Backbone Link
Zheng et al. [61] CIT LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Pythia-410M [174] Backbone Link
TGL [62] CPT GPT-2 base, GPT-2 large, GPT-Neo Backbone -
LOIRE [63] CPT GPT-2 small, GPT-2 large, BERT-base Backbone -
AdapterCL [64] CToD GPT-2 PEFT Link
SAPT [65] CIT T5-large, T5-XL, T5-XXL, LLaMA-2-7B, LLaMA-2-13B PEFT Link
MoRAL [66] CPT TinyLlama-1.1B [175], Phi-2-2.7B [176], LLaMA-2-7B PEFT -
I-LoRA [67] CIT LLaMA-2-7B PEFT Link
SLIM [68] CIT OpenChat-8B [177] PEFT -
P-Prompts [69] CIT BERT-base, T5-large PEFT Link
Q-Tuning [70] CIT BERT-base, T5-large PEFT -
WAVE-CRE [71] CRE BERT-base PEFT Link
O-LoRA [72] CIT T5-large, LLaMA-7B PEFT Link
DAS [73] CPT RoBERTa-base [178] Backbone Link
ARPER [74] CToD SCLSTM [179] Backbone Link
SEEKR [75] CIT LLaMA-2-7B-Chat, Vicuna-7B-v1.5 [180], Vicuna-13B-v1.5 Backbone Link
TaSL [76] CToD T5-small, T5-base, T5-large, LLaMA-7B Backbone & PEFT Link
DYNAINST [77] CIT BART-base [181] Backbone -
DaCL [78] CRE BERT-base Backbone Link
FVG [79] CIT LLaMA-2-7B-Chat, LLaMA-2-13B-Chat, LLaMA-3-8B-Chat, Mistral-7B-instruct-v2 [182] PEFT -
CT0 [80] CIT T5-small, T5-3B Backbone Link
Huang et al. [81] CIT OPT-125M [183], OPT-350M, OPT-1.3B, OPT-2.7B, OPT-6.7B, OPT-13B Backbone -
Anh et al. [82] CRE BERT-base, LLM2Vec [184] Backbone -
InsCL [83] CIT LLaMA-7B Backbone Link
HESIT [84] CToD GPT-2 Backbone -
Sun et al. [85] CIT Qwen2-0.5B [164], Qwen2-1.5B, LLaMA-2-7B, LLaMA-2-13B, Mistral-7B [182] Backbone Link
D-CPT Law [86] CPT Qwen1.5-0.5B [185], Qwen1.5-1.8B, Qwen1.5-4B Backbone -
LAMOL [87] CIT GPT-2 small Backbone Link
HMI-LAMOL [88] CIT GPT-2 small, BERT-base Backbone Link
PCLL [89] CToD GPT-2 small PEFT Link
DCL [90] CToD GPT-2 small PEFT -
SSR [91] CIT LLaMA-2-7B, LLaMA-2-7B-Chat, Alpaca-7B [186] Backbone Link
3.2 Architecture-based Approach
3.2.1 Backbone Expansion
Backbone expansion methods dynamically adapt network architectures to integrate new information by modifying parameter sparsity, expanding model capacity, or introducing new learning modules, without requiring full retraining. As shown in Figure 4 (a), these methods typically restructure internal components to balance the retention of previous knowledge with the acquisition of new capabilities. For instance, TPEM [60] proposes a continual learning framework for task-oriented dialogue systems based on iterative pruning, expansion, and masking. It begins by pruning redundant weights and retraining the retained parameters to achieve sparsity without sacrificing performance. To support new tasks, TPEM dynamically expands hidden dimensions based on pruning ratios and data volume. By analyzing early-stage optimization dynamics, Zheng et al. [61] reveal that performance drops often stem from disrupted alignment rather than erased knowledge. A layer-freezing strategy is proposed to stabilize low-level representations, yielding consistent improvements across multiple CL benchmarks. TGL [62] proposes a gradient-probing approach to localize time-sensitive entity knowledge in LLMs using salient span prediction tasks. Compared to generic continual pretraining or knowledge editing techniques, this method emphasizes the locality of knowledge in model parameters, demonstrating that selective layer updates can enable more efficient knowledge adaptation and mitigate catastrophic forgetting. LOIRE [63] introduces a plug-in layer expansion strategy with residual connections to ensure strict function preservation during depth-wise scaling. It further optimizes model growth across hidden dimensions, feed-forward layers, and attention heads. To address catastrophic forgetting, LOIRE employs an iterative distillation strategy, where intermediate models alternately act as teachers and students, enabling stable knowledge retention across evolving data distributions.
3.2.2 LoRA Expansion
To enable effective knowledge acquisition while preserving the backbone network’s original capabilities, parameter-efficient fine-tuning methods like LoRA and Adapter [187] have been widely adopted. These approaches leverage task-specific low-rank updates to the weight matrices, allowing the model to adapt to new tasks while maintaining stable performance on previously learned ones, thereby mitigating catastrophic interference. For example, AdapterCL [64] utilizes residual adapters to separate task-specific knowledge and adopts a perplexity-based inference strategy for adapter selection. SAPT [65] introduces a framework that aligns the learning and selection of PET blocks via a Shared Attentive Learning and Selection module to tackle catastrophic forgetting and knowledge transfer. It employs an Attentive Reflection Module to leverage generated pseudo samples for recalling shared attention from prior tasks, ensuring robust performance without requiring task identifiers during testing. MoRAL [66] integrates LoRA with mixture-of-experts routing to enable parameter-efficient adaptation and dynamic expert selection. I-LoRA [67] builds a dual-memory replay framework with a fast learner for quick adaptation and a slow learner for long-term knowledge consolidation. SLIM [68] introduces a dynamic mechanism that selectively activates or bypasses LoRA modules based on input characteristics. Specifically, it employs a learnable routing layer and sliding clustering to distinguish in-domain and out-of-domain inputs, directing general tasks through identity mappings and downstream tasks through LoRA adapters.
3.2.3 Prompt Expansion
As an alternative parameter-efficient fine-tuning approach for LLMs, prompt-based methods achieve continual learning by modularizing knowledge into prompts [188, 152], enabling efficient task adaptation and selective updates while reducing forgetting and interference. P-Prompts [69] learn a distinct prompt for each task and progressively concatenate them before feeding into a frozen backbone. It prevents forgetting by keeping task-specific prompts fixed after training while enabling forward transfer through the reuse of previous prompts. Q-Tuning [70] proposes a continual prompt tuning framework that maintains a prompt queue for old tasks and introduces new prompts for incoming tasks. It applies a low-rank adaptive knowledge aggregation mechanism for relevance reweighting, employs a PCA-based eviction rule to maintain constant queue size, and incorporates a globally shared prefix prompt with memory retention regularization to mitigate information loss across tasks. WAVE-CRE [71] proposes a continual relation extraction method that combines task-specific prompt pools with generative consolidation. The approach introduces a prompt pool for each task to capture intra-task variations and leverages a generative model to consolidate prior knowledge within shared parameters, effectively mitigating catastrophic forgetting without explicit data storage.
3.3 Regularization-based Approach
Regularization-based approaches, as illustrated in Figure 4 (b), mitigate catastrophic forgetting by enforcing training-time constraints that preserve critical parameters or feature representations from prior tasks. For example, O-LoRA [72] employs task-specific LoRA modules with orthogonal regularization, where each new task receives dedicated low-rank adaptation parameters. By enforcing orthogonality constraints on the LoRA updates, it effectively minimizes interference between task-specific parameter spaces, thereby substantially mitigating catastrophic forgetting. DAS [73] introduces a continual domain-adaptive pre-training framework for language models, employing a soft-masking mechanism to regulate parameter updates based on the importance of units for general or domain-specific knowledge. It preserves general knowledge through a robustness-based proxy and facilitates cross-domain transfer by contrasting learned knowledge with complete knowledge, thereby mitigating catastrophic forgetting and enabling knowledge integration without requiring domain IDs during fine-tuning. ARPER [74] applies an adaptive elastic weight consolidation regularization [30] to dynamically constrain important parameters based on vocabulary shifts across tasks, assigning stronger penalties to parameters critical for previous knowledge. SEEKR [75] employs knowledge distillation [158] on attention head outputs to preserve critical knowledge, while introducing task sensitivity and forgettability measures to dynamically identify and prioritize the most informative attention heads for efficient knowledge retention. TaSL [76] employs group-wise gradient-based importance scoring to identify task-specific and shared parameters, followed by fine-grained model averaging that selectively integrates knowledge from previous tasks while preserving critical parameters to mitigate forgetting and enhance bidirectional knowledge transfer. DYNAINST [77] is a hybrid method for lifelong in-context instruction learning that combines wide local minima regularization with dynamic instruction replay. Specifically, it introduces a KL-divergence term that encourages the model’s output distribution to remain smooth and less overconfident, thereby improving robustness and generalization to future tasks. NaCL [78] presents a noise-resistant contrastive learning framework for continual relation extraction, employing a rebooted auxiliary model to identify clean samples and performing noise-guided adversarial attacks to align noisy samples with their labels in feature space, enriching contrastive representations and mitigating forgetting. FVG [79] proposes function vector-guided training [189] that stabilizes function vectors through activation consistency and KL-divergence alignment. It reveals that forgetting primarily arises from biased function activations rather than parameter overwriting, positioning function vectors as effective tools for diagnosing forgetting.
3.4 Replay-based Approach
3.4.1 Direct Replay
An intuitive replay-based approach involves retaining selected data samples from previous tasks or their compressed representations. As shown in Figure 4 (c), these stored exemplars are strategically reintroduced during the training of new tasks to prevent catastrophic forgetting. For example, CT0 [80] introduces a rehearsal mechanism by replaying 1% of historical data from each task, which allows the model to continually adapt to new tasks while preserving its previously acquired instruction-following ability. Huang et al. [81] propose a tool-augmented mechanism combined with a small replay buffer, which helps improve the model’s adaptation to new tasks while effectively mitigating the forgetting of old tasks, thus enhancing overall performance across both seen and unseen tasks. Anh et al. [82] combine old and new task samples into complex dual-label texts to mitigate data scarcity. They enhance model robustness through adversarial-like perturbations and employ specialized loss functions to consolidate relation representations while reducing overfitting. The method further improves generalization by selectively applying Sharpness-Aware Minimization [190] to current-task training data.
Moreover, several methods adopt dynamic data replay strategies. InsCL [83] adjusts the replay ratio of historical data dynamically based on task similarity. It also proposes a new metric for measuring instruction informativeness, favoring high-quality samples with richer information for training. HESIT [84] tracks how much each sample changes the model early in training (i.e. the gradient of model parameters with respect to the sample’s weight), then picks and replays the most impactful examples to effectively reduce forgetting of old tasks. Sun et al. [85] demonstrate that performance degradation in continual learning stems from the original prompts’ inability to effectively steer the model’s rationale generation. They introduce the Rationale-Guidance Difficulty metric to measure how hard it is for a prompt to guide reasoning, and they use this metric to replay old data as needed to keep and restore past task performance. There are also works that aim to optimize the allocation ratio between replay data and current task data in continual learning. For example, D-CPT Law [86] establishes a scaling law to optimize the mixture ratio between general and domain-specific corpora. This method extends the Chinchilla scaling law [191] by incorporating mixture ratio as a parameter and introducing a Domain-specific Learnable Coefficient to enable cross-domain generalization in limited-data scenarios.
3.4.2 Generative Replay
Generative replay methods leverage LLMs to synthesize pseudo-samples from previous tasks, enabling memory consolidation without needing to store real historical data. For example, LAMOL [87] introduces a generation token to guide the model in producing pseudo-samples alongside new task learning. This approach enhances the model’s ability to retain old knowledge without incurring extra computational costs. HMI-LAMOL [88] proposes a brain-inspired generative replay mechanism, HMI, which uses BERT to encode training samples into compressed “memory traces”. These traces condition the generation of old task data, effectively alleviating sample imbalance and quality degradation, thus mitigating catastrophic forgetting in language models. Furthermore, PCLL [89] constructs natural language prompts for each task and employs a conditional variational autoencoder to model the task-specific latent distribution, generating high-quality pseudo-samples for replay. DCL [90] replaces the Gaussian distribution used in PCLL with a Dirichlet distribution to model the latent variables, enabling better representation of structural differences across tasks, and introduces Jensen–Shannon knowledge distillation at the logit layer to improve the model’s retention of old knowledge. SSR [91] further introduces a self-synthesizing reproduction framework, which utilizes the context-learning ability of a base LLM to generate synthetic samples. These samples are refined by the latest fine-tuned model, and high-quality ones are selected via clustering algorithms, enabling the retention of prior knowledge without requiring historical data or additional generative models.
3.5 Discussions
4 Continual Learning in Multimodal Large Language Models
Recent Multimodal Large Language Models [192], built upon advanced Large Language Models, have demonstrated impressive visual understanding capabilities by integrating visual encoders, cross-modal projectors, and leveraging large-scale image-text corpus during training. Continual learning in MLLMs seeks to equip them with task-specific abilities across domains such as healthcare [193, 194], finance [195, 196], and autonomous driving [197, 198], without forgetting their previously learned knowledge. In contrast to discriminative models that predict output through classifiers, MLLMs generate responses by next-token prediction in an autoregressive way, making most continual learning methods designed for discriminative models less applicable. Additionally, compared to LLMs, MLLMs need to simultaneously process information from multiple modalities [199], thus posing a greater challenge.
4.1 Problem Setup
4.1.1 Training objectives
The visual instruction tuning dataset $\mathcal{D}=\{(\textbf{X}_{v}^{j},\textbf{X}_{ins}^{j},\textbf{X}_{a}^{j})_{j=1}^{N}\}$ for fine-tuning a MLLM $f_{\boldsymbol{\theta}}$ typically consists of image input $\textbf{X}_{v}$, user instruction $\textbf{X}_{ins}$ and target answer $\textbf{X}_{a}$, where $N$ represents the total number of training samples. Given an image-instruction pair with a response of length $L$, the MLLM predicts the next token autoregressively based on all preceding tokens:
$p(\textbf{X}_{a}|\textbf{X}_{v},\textbf{X}_{ins})=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(x_{i}|\textbf{X}_{v},\textbf{X}_{ins},\textbf{X}_{a,<i}),$ (15)
where $\theta$ denotes the trainable parameters during fine-tuning, $\textbf{X}_{a,<i}$ are the response tokens from all previous turns before the current prediction token $x_{i}$. In the continual learning framework for MLLMs, the optimization objective and loss function maintain the same formulation as the LLM case in Eq.(13) and Eq.(14), with the distinction that the input-output pairs now involve multimodal data. Here, $\boldsymbol{x}$ and $\boldsymbol{y}$ represent the multimodal inputs $(\textbf{X}_{v},\textbf{X}_{ins})$ and the target answer $\textbf{X}_{a}$, respectively.
4.1.2 Backbone
To handle inputs from multiple modalities, MLLMs typically begin by extracting modality-specific features using dedicated encoders. These features are subsequently concatenated and fed into a pre-trained language model, which generates responses in an autoregressive manner. Based on existing literature, we summarize the backbone architectures commonly used in continual learning methods for MLLMs as follows.
  • VL-T5 [118] employs two distinct visual encoders for different modalities: Faster R-CNN [8] for extracting region features in visual question answering, and 3D ResNeXt-101 [200] for capturing video motion features in video question answering. The language model in VL-T5 comprises 12 encoder and decoder blocks, each containing a multi-head attention layer with 12 heads, and the embedding dimension is set as 768.
  • InstructBLIP [201] utilizes a Query Transformer (Q-Former) to extract visual features from a frozen image encoder (ViT-G/14 [202]). The input to the Q-Former contains a set of $K$ learnable query embeddings, which interact with the image encoder’s output through cross-attention. The Q-Former outputs are then passed through a linear projection before being fed into the frozen LLM. For the LLM, existing methods [92, 104] typically employ FlanT5-XL which is an instruction-tuned model based on the encoder-decoder Transformer T5 [168]. During training, only the projection layer of the current task in the Q-Former and its corresponding key are trainable.
  • LLaVA [15, 18] series has emerged as one of the most prominent MLLMs, distinguished by its impressive visual understanding capabilities and streamlined architecture. It employs Vicuna [203] as LLM and pre-trained CLIP visual encoder ViT-L/14@336 [204] to extract visual embeddings from images. To align visual and textual modalities, LLaVA employs a simple yet effective linear projection layer to project visual features into a unified representation space.
In addition to the MLLM backbones mentioned above, some methods incorporate customized architectural designs tailored to their specific settings (e.g., Qwen-VL series [17, 19], MiniGPT-v2 [16], InternVL [205], etc.), which are detailed in Table II.
Figure 5: Three types of continual learning approaches in Multimodal Large Language Models.
4.1.3 Benchmark
To evaluate and reveal the continual learning capabilities of MLLMs, several benchmarks have been developed. In this part, we focus on a selection of the most representative benchmarks to provide a detailed discussion. For details regarding the benchmarks and datasets employed by each method, we refer readers to their original papers.
  • CLiMB [206] is introduced to investigate continual learning in vision-and-language settings, supporting both multimodal and unimodal tasks. It consists of two phases: (1) Upstream Continual Learning of Multimodal Tasks: The pre-trained multimodal model is trained on a sequence of vision-and-language tasks and evaluated after each task to assess Forgetting of previous tasks and Knowledge Transfer to subsequent tasks. (2) Downstream Transfer to Low-Shot Tasks: After each multimodal task, the model is evaluated for its Low-Shot Transfer capability on both multimodal and unimodal downstream tasks. The experiments in CLiMB demonstrate that existing continual learning strategies struggle to generalize across sequences of multimodal tasks and fail to enable effective low-shot adaptation to downstream multimodal or unimodal tasks.
  • VQACL [118] designs dual-level task sequence using two datasets: VQA2.0 [207] and NExT-QA [208]. The outer level consists of $T$ linguistic-driven tasks, where each task corresponds to learning a specific reasoning skill. For example, for the ’Count’ task, its training data mainly contain the examples that can teach the model how to count, such as ‘How many’ and ‘What number is’. At the inner level, each linguistic-driven task is further composed of a sequence of visual-driven subtasks, with each subtask containing images from a specific object group. Furthermore, VQACL focuses on evaluating the compositionality of the model’s reasoning skills (e.g., counting, color) and visual concepts (e.g., cat, surfboard). Specifically, a visual-driven subtask is randomly removed from each linguistic-driven task during training and used as a new composition for testing.
  • CoIN [94] consists of carefully designed datasets covering 8 tasks, including general image question answering, visual reasoning, knowledge-grounded image question answering, etc. Additionally, CoIN proposes to evaluate MLLMs from two perspectives: (1) Truth Alignment: The ability to generate the correct result in the desired format, adhering to task instructions, is a fundamental requirement for instruction tuning. (2) Reasoning Capability. The performance of MLLMs depends not only on their ability to follow instructions but also on the knowledge they retain and integrate over time.
  • MLLM-CL [107] introduces two settings named domain continual learning (DCL) and ability continual learning (ACL), providing a comprehensive and realistic evaluation for continual learning of MLLMs. Specifically, DCL focuses on learning domain-specific knowledge continually on several mainstream domains (remote sensing, medical, autonomous driving, science, and finance), and evaluated on independently and identically distributed (IID) test sets. Differently, ACL aims to incorporate fundamental abilities (OCR, math & logic, visual perception and, GUI agent), evaluating non-IID scenarios.
4.2 Architecture-based Approach
4.2.1 Backbone Expansion
As shown in Figure 5 (a), backbone expansion methods dynamically extend model architectures through modular components, enabling adaptive knowledge acquisition for diverse tasks. Eproj [92] establishes the first continual instruction tuning benchmark for MLLM. To address catastrophic forgetting, the framework dynamically selects optimal Q-Former projection layers through multimodal similarity scoring between input samples and preserved training features, while integrating adaptive regularization constraints with a strategic old-sample replay mechanism to maintain stable model performance. TAM-CL [93] proposes a novel transformer-based continual learning framework that dynamically expands model capacity through task-specific attention layers with learnable tokens. The approach simultaneously preserves acquired knowledge through intermediate teacher-student distillation while employing a compact memory buffer for selective experience replay, ensuring robust performance across sequential tasks.
TABLE II: Summary of continual learning methods for Multimodal Large Language Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Benchmark: Composite benchmark refer to method that combine multiple datasets without a unified name. Tag: I = Image, T = Text, V = Video, A = Audio.
Arch Reg Rep
Eproj [92] Composite InstrutBLIP, BILP2 [209] I T Backbone -
TAM-CL [93] Composite ViLT [210] I T Backbone Link
MoELoRA [94] CoIN LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen-VL-Chat, MiniGPT-v2 I T PEFT Link
Continual LLaVA [95] COAST LLaVA-1.5-7B I T PEFT Link
HiDe-LLaVA [96] CoIN, UCIT LLaVA-1.5-7B, InternVL-Chat-7B I T PEFT Link
LLaCA [97] CoIN LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen-VL-Chat I T PEFT Link
SMoLoRA [98] CVIT LLaVA-1.5-7B I T PEFT -
Cl-MoE [99] VQACL VL-T5, LLaVA-1.5-7B I T PEFT Link
LLaVA-CMoE [100] CoIN LLaVA-1.5-7B I T PEFT -
DISCO [101] FCIT LLaVA-1.5-7B I T PEFT -
TRIPLET [102] Composite ALBEF [211], FLAVA [212] I T PEFT -
ModalPrompt [103] CoIN LLaVA-1.5-7B I T PEFT Link
Fwd-Prompt [104] Composite InstructBLIP, BLIP2 I T PEFT -
ColPro [105] Composite LLaMA-7B, CLIP ViT-L/14 V T PEFT Link
CluMo [106] CLOVE ALBEF I T PEFT Link
MR-LoRA [107] MLLM-CL LLaVA-1.5-7B I T PEFT Link
SCD [108] MDL-VQA ViLT I T Backbone -
CS-VQLA [109] Composite ResNet [3], VisualBERT [213] I T Backbone Link
Chen et al [110] Composite InstructBLIP I T Backbone -
CrossSDC [111] Composite iQuery [214] V A Backbone Link
MAFED [112] Composite UNITER [215], ViLT, VL-Pythia I T Backbone Link
[…]
QUAD [114] VQACL VL-T5 I T Backbone Link
Model Tailor [115] Composite InstructBLIP, LLaVA-1.5-7B I T Backbone & PEFT Link
SPIDER [116] Composite VILA-1.5-3B, LLaVA-1.5-7B I T Backbone -
SEFE [117] CoIN-ASD LLaVA-1.5-7B I T PEFT Link
VQACL [118] VQACL VL-T5 I T Backbone Link
SGP [119] CLOVE UniVQA I T Backbone Link
ProtoGroup [120] VQACL VL-T5 I T Backbone -
Lin et al [121] Composite EM-VLM4AD [217] I T Backbone -
Adapt-$\infty$ [122] Composite LLaVA-1.5-7B I T PEFT Link
OASIS [123] MICVIT LLaVA-1.5-7B, Qwen-VL-2.5-7B I T PEFT -
4.2.2 LoRA Expansion
[…]

### v4

@@ 4.2.1. Backbone Expansion
QUAD (Marouf et al., 2025) VQACL VL-T5 I T Backbone Link
Model Tailor (Zhu et al., 2024c) Composite InstructBLIP, LLaVA-1.5-7B I T Backbone & PEFT Link
SPIDER (Huang et al., 2024b) Composite VILA-1.5-3B, LLaVA-1.5-7B I T Backbone -
SEFE (Chen et al., 2025a) CoIN-ASD LLaVA-1.5-7B I T PEFT Link
LoRASculpt (Liang et al., 2025) Composite LLaVA-1.5-7B I T PEFT Link
LLaVA-c (Liu et al., 2025c) Composite LLaVA-1.5-7B I T PEFT -
VQACL (Zhang et al., 2023d) VQACL VL-T5 I T Backbone Link
SGP (Lei et al., 2023) CLOVE UniVQA I T Backbone Link
ProtoGroup (Zhang et al., 2025a) VQACL VL-T5 I T Backbone -
Lin et al (Lin et al., 2025) Composite EM-VLM4AD I T Backbone -
Adapt-$\infty$ (Maharana et al., 2024) Composite LLaVA-1.5-7B I T PEFT Link
OASIS (Lee et al., 2025) MICVIT LLaVA-1.5-7B, Qwen-VL-2.5-7B I T PEFT -
4.2.2. LoRA Expansion
LoRA expansion methods enhance MLLMs by adding lightweight low-rank adapters to a fixed backbone, balancing learning flexibility and memory stability for efficient continual learning. MoELoRA (Chen et al., 2024f) embeds LoRA modules as experts within a mixture of experts, using a learnable router for dynamic expert selection, evaluated on the CoIN benchmark. Continual LLaVA (Cao et al., 2024) introduces COAST, a benchmark assessing domain shifts, capability growth, and dataset evolution, mitigating forgetting via a dual-embedding combining instruction-aware and cross-task contextual features alongside selective LoRA modules. HiDe-LLaVA (Guo et al., 2025b) decomposes the model via layer-wise CKA similarity into task-specific expansions and general fusion, guided by dual-modality prototype matching for adaptive LoRA selection, with parameter-efficient merging to preserve shared knowledge, supported by the UCIT benchmark. MR-LoRA (Zhao et al., 2025b) employs architectural decoupling to preserve plasticity and uses multimodal routing based on cross-modal representations for dynamic expert selection in complex scenarios. LLaCA (Qiao et al., 2024) balances stability and plasticity with gradient-guided EMA weight adaptation and Taylor expansion, enhanced by semantic instruction grouping for robust anti-forgetting. SMoLoRA (Wang et al., 2024a) introduces a separable mixture of low-rank adapters with distinct routing for vision and language tasks, dynamically fusing LoRA blocks to reduce dual forgetting, evaluated on the CVIT benchmark. CL-MoE (Huai et al., 2025) proposes dual-router MoE combining instance- and task-level routing, plus momentum-based parameter updates for shared and task-specific experts to balance retention and acquisition. LLaVA-CMoE (Zhao et al., 2025a) dynamically expands experts based on activation frequency, leveraging hierarchical routing and VAE-modeled task distributions to avoid interference and preserve zero-shot capabilities. FCIT (Guo et al., 2025c) presents DISCO for federated continual instruction tuning, dynamically managing task subspaces via identity token matching and selective activation, establishing the first FCIT benchmark addressing data heterogeneity and continual learning challenges. MR-LoRA (Zhao et al., 2025b) combines domain-specific LoRA modules with an MLLM-based routing mechanism to dynamically select optimal experts, preventing catastrophic forgetting via parameter isolation and few-shot adaptation. D-MoLE (Ge et al., 2025) introduces dynamic layer-wise LoRA expert allocation and gradient-based inter-modal curriculum to optimize continual multimodal instruction tuning, achieving adaptive parameter distribution across transformer layers while balancing modality-specific updates through task-aware difficulty assessment. BranchLoRA (Zhang et al., 2025b) uses a shared matrix $A$ in LoRA modules to capture task-invariant patterns, alongside task-specific matrices $B$ equipped with flexible tuning–freezing mechanisms and task-specific routers, thereby effectively mitigating catastrophic forgetting. ProgLoRA (Yu et al., 2025a) employs a progressive LoRA pool with task-aware allocation and recall to mitigate forgetting and enhance knowledge transfer in multimodal continual instruction tuning.
4.2.3. Prompt Expansion
Prompt expansions are a standard technique used to improve the continual learning capability of MLLMs. For example, TRIPLET (Qian et al., 2023) proposes the first multimodal prompt learning framework for continual visual question answering, introducing decoupled prompts across modalities, transformer layers, and tasks, combined with query-matching and cross-modal interactions to model complex dependencies and mitigate forgetting without rehearsal buffers. ModalPrompt (Zeng et al., 2024b) presents a dual-modality guided prompt learning approach that constructs task-specific prototype prompts to preserve prior knowledge without replay, using image-text guided prompt selection to dynamically match relevant prompts at inference. Fwd-Prompt (Zheng et al., 2024) applies prompt tuning with gradient projection, allocating orthogonal subspaces to prevent interference by projecting gradients to residual spaces for old knowledge retention while reusing core pre-trained knowledge to boost forward transfer. ColPro (Cai et al., 2024) introduces collaborative prompting by integrating task-specific question constraints, knowledge acquisition, and visual temporal awareness prompts into a pre-trained LLM, leveraging complementary learning with General and Expert prompts, negative guiding, autoregressive temporal dynamics, and multimodal distillation to enhance video QA and reduce forgetting. CluMo (Cai and Rostami, 2024) proposes a cluster-based modality fusion prompting method with task-specific visual and textual prompt keys learned via K-means clustering, selecting optimal prompts from a shared pool using a two-stage training and knowledge distillation strategy to improve generalization and mitigate forgetting while keeping the backbone frozen.
4.3. Regularization-based Approach
4.3.1. Feature-Level Regularization
To overcome catastrophic forgetting, the idea of Knowledge Distillation has also demonstrated strong potential in continual learning approaches for MLLM. One line of work uses current task data directly for distillation, aiming to retain previous knowledge without relying on past data. For example, SCD (Lao et al., 2023) distills instance- and domain-relevant knowledge by comparing teacher responses to original and counterfactual samples, employing metric learning to avoid harmful domain-specific features while acquiring transferable knowledge. CS-VQLA (Bai et al., 2023b) introduces rigidity-plasticity-aware and self-calibrated heterogeneous distillation to balance retention and adaptability in robotic surgery tasks. Chen et al. (Chen et al., 2024b) propose an LLM-assisted distillation loss that leverages LLMs’ generalization to handle domain shifts and data imbalance. CrossSDC (Pian et al., 2024b) preserves cross-modal semantic similarity over time and can be integrated into various frameworks to prevent forgetting in continual audio-visual sound separation.
Another approach uses distillation with past task data or generated pseudo-samples. MAFED (Nikandrou et al., 2024) mixes a small portion of old samples into new task data and applies modality-aware feature distillation, weighting questions and visual tokens by importance. MoInCL (Pian et al., 2024a) proposes a pseudo-target generation module to synthesize input-target pairs for various task types. These pseudo samples are then mixed with current task data for knowledge distillation, effectively mitigating catastrophic forgetting. QUAD (Marouf et al., 2025) uses only past task questions for regularization to avoid overfitting and introduces attention consistency distillation to preserve intra- and inter-modal attention across tasks.
4.3.2. Parameter-Level Regularization
[…]